# Relative Camera Pose Regression

### v1

@@ 4.1 Dataset Construction
We leverage the co-visibility estimation pipeline introduced in RUBIK [26] to generate pixel-level co-visibility annotations for image pairs. In brief, this pipeline uses monocular metric depth predictions from UniDepth [29] and surface normals from Depth Anything V2 [48], combined with camera poses from COLMAP [32] reconstructions, to automatically classify each pixel as either co-visible, occluded, or outside FOV with respect to another image.
Starting from the nuScenes training set, we apply this pipeline to all possible image pairs within each scene, resulting in approximately 34 million annotated pairs. For our experiments, we created two dataset variants, each containing 2.5 million image pairs:
  • Cub3-50: Image pairs with at least 50% overlap, similar to the criterion used in CroCo [46]. This dataset provides a direct comparison point with existing cross-view completion approaches.
@@
Fig. 3 shows examples or our co-visibility annotations on image pairs from Cub3, illustrating how our process classifies pixels into the three categories across varying levels of difficulty.
It is also worth noting that our annotation process inherits some limitations from the underlying monocular depth estimation models. The depth predictions may occasionally struggle with challenging scenarios such as reflective surfaces, transparent objects, or regions with complex geometry or far-away objects. Consequently, some annotations in our dataset, particularly the distinction between co-visible and occluded pixels, may contain noise. Despite these imperfections, our experiments show that the scale and diversity of Cub3 enable effective training of robust co-visibility segmentation models.
@@ 4.2 Dataset Statistics
Figure 3: Co-visibility annotation examples from Cub3. For each image pair, we show the corresponding co-visibility maps with color-coding for co-visible, occluded, and outside FOV regions. Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes. Let us highlight that some annotations, particularly the distinction between co-visible and occluded pixels, may contain noise, see text for details.
Figure 4: Distributions of overlap, scale ratio, and viewpoint angle in Cub3-all and Cub3-50.
Fig. 4 illustrates the distribution of our datasets across three geometric criteria introduced in RUBIK: overlap percentage, scale ratio, and viewpoint angle. The Cub3-all dataset’s distribution is heavily skewed towards very challenging pairs, which are particularly difficult for cross-completion methods like CroCo. In contrast, our co-visibility segmentation approach is designed to benefit from such challenging cases effectively.
Given our computational resources, Cub3 is currently limited to autonomous driving scenarios from nuScenes. But such scenarios have very important applications, and given its size, variety, and challenges, Cub3 is already sufficient to validate Alligat0R—very much like the initial version of CroCo [45] was demonstrated as a ”proof-of-concept”:
  • Scale: With 2.5 million annotated pairs, Cub3 provides sufficient data to train and validate our approach.
  • Geometric variety: It covers a wide range of overlaps, scales and viewpoint angles within the driving context.
@@
  • Application relevance: Autonomous driving represents an important application domain with significant research interest.
Our experiments on the RUBIK benchmark, which was specifically designed to evaluate challenging autonomous driving scenarios, demonstrate the effectiveness of our approach. The promising results we obtain motivate future work to extend Cub3 to more diverse environments beyond urban driving scenes. However, even within this specific domain, our method shows significant improvements over CroCo, particularly in geometrically challenging configurations.
@@ 5.1 Implementation Details
We implement Alligat0R using PyTorch and conduct all experiments on NVIDIA A100 GPUs. The model architecture follows the design described in Sec. 3.1, utilizing a ViT-based encoder and transformer decoder backbone similar to CroCo, with 24 layers for the encoder and 12 for the decoder. The input images are resized to have a maximum dimension of 512 pixels while maintaining their aspect ratio, resulting in an input resolution suitable for capturing sufficient detail while remaining computationally efficient. For pre-training, we use the AdamW optimizer with a learning rate of 1.5e-4, weight decay of 0.05, and a batch size of 32 per GPU. We employ a cosine learning rate schedule with 2 epochs of warmup and train for 25 epochs on our Cub3 datasets.
For fine-tuning on relative pose regression, we follow the two-phase approach described in Sec. 3.3. In the first phase, we freeze the backbone and train only the pose regression head for 5 epochs with a learning rate of 1e-4. In the second phase, we unfreeze the entire network and jointly train with both the pose regression loss and co-visibility segmentation loss for an additional 10 epochs with a learning rate of 5e-5.
The comprehensive architecture of Alligat0R is illustrated in Fig. 2, showing both the pre-training and fine-tuning phases. During pre-training, the model learns to segment pixels in each view as co-visible, occluded, or outside the field of view with respect to the other view. During fine-tuning, we introduce a pose regression head while maintaining the segmentation capability to leverage the geometric understanding acquired during pre-training.
@@ 5.2 Main Results
Table 1: Results on RUBIK [26] Benchmark for Relative Pose Regression. See text for details.
Pre-training Training set Backbone Finetune set $5^{\circ}$ / 2m $10^{\circ}$ / 5m AUC @ 5/10/20
CroCo Cub3-50 [Uncaptioned image] Cub3-50 19.10 28.80 15.6 / 24.7 / 33.9
CroCo Cub3-all [Uncaptioned image] Cub3-all 8.90 25.20 8.5 / 22.0 / 40.8
Alligat0R Cub3-50 [Uncaptioned image] Cub3-50 24.40 32.10 14.8 / 23.1 / 31.2
Alligat0R Cub3-all [Uncaptioned image] Cub3-all 55.30 82.30 31.1 / 54.0 / 71.8
Alligat0R Cub3-all [Uncaptioned image] Cub3-all 60.30 81.90 41.6 / 61.2 / 75.5
From scratch - [Uncaptioned image] Cub3-all 29.10 52.30 27.1 / 43.7 / 58.7
@@
Relative Pose Regression Performance. We evaluate our proposed Alligat0R pre-training approach on the relative pose regression task and compare it with CroCo pre-training using the same architecture and fine-tuning protocol. Tab. 1 presents the results on the RUBIK [26] benchmark, which is specifically designed to evaluate performance across different geometric challenges in autonomous driving scenarios.
@@
As shown in Tab. 1, Alligat0R consistently outperforms CroCo across different training and evaluation configurations. Notably, when pre-trained on Cub3-all and fine-tuned with the backbone frozen, Alligat0R achieves significantly higher performance and reaches state-of-the-art on the RUBIK benchmark (55.30% success rate at $5^{\circ}$/2m and 82.30% at $10^{\circ}$/5m) compared to the CroCo model pre-trained on the same dataset (which only achieves 8.90% and 25.20% respectively in the frozen backbone configuration), confirming the results in the original CroCo paper. This substantial improvement demonstrates the effectiveness of our co-visibility segmentation pre-training approach, especially for challenging scenarios with varying degrees of overlap.
Impact of Training Data Distribution. Our results reveal an interesting pattern regarding the impact of training data distribution. When trained on pairs with at least 50% overlap, CroCo achieves modest performance (19.10% accuracy at $5^{\circ}$/2m and 28.80% at $10^{\circ}$/5m). However, when training on the full distribution, Alligat0R shows remarkable improvement (from 24.40% to 55.30% at $5^{\circ}$/2m and from 32.10% to 82.30% at $10^{\circ}$/5m). This confirms our hypothesis that explicitly modeling co-visibility through segmentation enables more effective learning from challenging image pairs with limited overlap.
@@
Figure 5: Performance of CroCo and Alligat0R across different geometric challenges. Results show accuracy at the $5^{\circ}$/2m threshold for models trained on different datasets (Cub3-50 or Cub3-all) and for frozen backbones. Alligat0R trained on Cub3-all consistently outperforms other configurations, particularly for challenging cases with low overlap, large scale differences, and extreme viewpoint changes.
@@
Fig. 5 further illustrates this advantage by breaking down performance across different overlap percentages, scale ratios and viewpoint angles. While both methods perform well on high-overlap pairs (80-100%), Alligat0R trained on Cub3-all maintains strong performance even as overlap decreases, whereas CroCo’s accuracy drops dramatically for pairs with less than 40% overlap. This demonstrates Alligat0R’s ability to generalize across a wide range of geometric configurations, which is crucial for real-world applications where overlap cannot be guaranteed.
@@
Fine-tuning Strategy Comparison. We also compare different fine-tuning strategies, specifically the impact of freezing versus unfreezing the backbone during fine-tuning. For Alligat0R, unfreezing the backbone during fine-tuning leads to even better performance and yielding the best results (60.30% at $5^{\circ}$/2m and 81.90% at $10^{\circ}$/5m compared to 55.30% and 82.30% with the frozen backbone), suggesting that our pre-training approach learns transferable features that can be effectively adapted to the pose regression task.
@@
Comparison with Training from Scratch. To isolate the impact of our pre-training approach, we compare the performance of a model trained from scratch for relative pose regression versus those initialized with Alligat0R or CroCo pre-training. The final row in Tab. 1 shows that training from scratch achieves notably lower performance (29.10% at $5^{\circ}$/2m and 52.30% at $10^{\circ}$/5m) compared to our pre-trained models trained on the same dataset (Cub3-all), highlighting the value of the representations learned during pre-training.
@@
Table 2: RUBIK Benchmark. The results are reported from [26]. Success rate (in %) for each method across individual geometric criterion bins. Best and second-best values for each column are shown in bold and underlined respectively.
@@
Detector-based methods
ALIKED+LightGlue [52] 53.4 95.8 68.2 38.0 12.7 62.0 31.0 13.1 1.6 50.6 46.0 28.3 2.0 36.8 45
DISK+LightGlue [39] 54.2 91.4 65.9 38.7 11.8 60.4 30.8 11.6 2.4 50.3 43.8 27.4 2.7 35.9 69
SP+LightGlue [6] 64.8 93.3 68.0 36.4 10.9 61.2 28.4 12.5 1.4 49.9 43.0 28.2 0.9 35.7 43
SIFT+LightGlue [27] 68.2 92.1 61.4 32.3 9.9 57.3 26.9 9.6 1.7 49.8 39.7 23.7 0.5 33.1 194
DeDoDe v2 [10] 89.8 93.3 54.8 26.7 7.9 60.4 16.3 3.2 0.9 49.3 35.4 19.9 0.3 30.4 282
XFeat [30] 85.4 67.4 24.3 5.2 0.9 32.1 2.4 0.1 0.0 34.4 8.3 7.1 0.0 14.2 54
XFeat* [30] 62.4 69.1 27.6 7.6 1.5 32.8 4.5 0.6 0.0 33.8 9.4 9.2 0.0 15.1 82
XFeat+LighterGlue [30] 64.6 91.7 59.1 26.2 8.1 56.6 20.9 4.6 0.2 48.0 33.4 21.4 1.2 30.1 43
@@
LoFTR [35] 87.2 88.4 47.2 17.5 5.0 51.6 10.1 2.3 0.6 43.2 27.9 15.1 0.0 24.9 185
ELoFTR [44] 56.4 90.3 50.8 22.1 6.3 51.2 15.6 4.4 0.7 42.2 30.8 18.2 0.1 26.6 124
ASpanFormer [4] 72.2 72.3 44.5 21.9 7.4 46.0 14.9 6.9 1.6 42.5 27.2 16.0 0.1 24.8 108
RoMa [11] 67.0 98.3 84.5 52.7 20.2 71.2 43.2 26.6 8.3 57.5 56.2 44.1 3.0 47.3 614
DUSt3R [43] 81.8 97.4 90.8 58.4 30.4 73.3 57.9 40.1 9.9 67.4 55.3 50.0 35.2 54.8 257
MASt3R [24] 52.0 97.5 89.6 61.0 28.4 71.2 52.3 42.5 13.8 53.5 65.6 54.5 14.1 53.6 173
@@ 5.3 Qualitative Analysis
Co-visibility and Reconstruction Visualization. Fig. 6 shows qualitative examples of the co-visibility segmentation and CroCo reconstructions produced by our models on pairs from RUBIK. Alligat0R successfully identifies co-visible regions, occluded areas, and regions outside FOV across varying degrees of overlap and viewpoint changes. These visualizations provide insights into the geometric understanding acquired by our model. Conversely, the best CroCo model succeeds in reconstructing masked regions for pairs with high overlaps, but completely fails on more difficult pairs.
Figure 6: Qualitative comparison between CroCo trained on Cub3-50 and Alligat0R. For each row from left to right: reference image, target image, masked target image, CroCo reconstruction, and Alligat0R segmentations (co-visible, occluded, outside FOV). CroCo struggles with reconstructing regions with limited overlap, while Alligat0R often correctly identifies co-visible, occluded, and outside FOV regions across varying degrees of overlap and viewpoint changes.
@@ 5.4 Comparison with state-of-the-art on RUBIK
In Tab. 2, we evaluate the performance of Alligat0R fine-tuned for metric relative pose estimation against state-of-the-art methods on RUBIK [26]. Let us highlight that:
@@
This gives an advantage to Alligat0R over the other methods. However, RUBIK allows the evaluation over large ranges of three geometric criteria—overlap, scale ratio and viewpoint angle—which we found interesting. Despite this advantage, we argue that Alligat0R’s performance remains remarkable, as pretraining with CroCo, in contrast, leads to a significant performance drop (see Tab. 1).
With an overall success rate of 60.3%, Alligat0R ranks first. Notably, it significantly outperforms all other methods on the hardest bin across the three geometric criteria, which confirms the advantage of its pretraining on difficult image pairs. While Alligat0R’s training set was highly skewed toward challenging pairs (see Fig. 4), one might expect a performance drop on easier bins. However, this does not happen: for instance, Alligat0R still ranks first for small scale ratios and small viewpoint angles. Finally, Alligat0R is among the fastest methods, as it directly regresses the relative pose rather than producing intermediate correspondences.
5.5 Future Directions
Our results demonstrate the effectiveness of co-visibility segmentation for pre-training. From this idea, we believe that several promising directions can be explored:
  1. 1.
    Segmentation Formulations: Investigating different ways to formulate the segmentation task, including binary vs. three-class approaches.
  2. 2.
    Curriculum Learning: Experimenting with curriculums that gradually increase the difficulty of training examples to see if this could further improve performance.
  3. 3.
    Multi-task Learning: Studying whether jointly training on multiple downstream tasks can leverage the geometric understanding provided by our co-visibility segmentation approach.
  4. 4.
    Diverse Environments: Extending Alligat0R to more diverse indoor and outdoor environments beyond autonomous driving scenarios.

### v2

@@ 4.1 Dataset Construction
nuScenes – We leverage the covisibility estimation pipeline introduced in RUBIK [loiseau2025rubik] to generate pixel-level covisibility annotations for image pairs from nuScenes. In brief, this pipeline uses monocular metric depth predictions from UniDepth [piccinelli2024unidepth] and surface normals from Depth Anything V2 [yang2024depth], combined with camera poses from COLMAP [schonberger2016structure] reconstructions, to automatically classify each pixel as either covisible, occluded, or outside FOV with respect to another image. We apply this pipeline to all possible image pairs within each scene, resulting in approximately 34 million annotated pairs.
ScanNet – The pipeline is similar but simpler for ScanNet, as we rely on the ground truth depths and camera poses to get all covisibility annotations.
For our experiments, we created two dataset variants, each containing 5M image pairs (2.5M from nuScenes and 2.5M from ScanNet):
  • Cub3-50: Image pairs with at least 50% overlap, similar to the criterion used in CroCo [weinzaepfel2023croco]. This dataset provides a direct comparison point with existing cross-view completion approaches.
@@
Figure˜3 shows examples or our covisibility annotations on image pairs from Cub3, illustrating how our process classifies pixels into the three categories across varying levels of difficulty.
It is also worth noting that our annotation process, especially on nuScenes, inherits some limitations from the underlying monocular depth estimation models. The depth predictions may occasionally struggle with challenging scenarios such as reflective surfaces, transparent objects, or regions with complex geometry or far-away objects. Consequently, some annotations in our dataset, particularly the distinction between covisible and occluded pixels, may contain noise. Despite these imperfections, our experiments show that the scale and diversity of Cub3 enable effective training of robust covisibility segmentation models.
@@ 4.2 Dataset Statistics
Figure 3: Covisibility annotation examples from Cub3 for nuScenes (top) and ScanNet (bottom). For each image pair, we show the corresponding covisibility maps with color-coding for covisible, occluded, and outside FOV regions. Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes. Let us highlight that some annotations, particularly the distinction between covisible and occluded pixels, may contain noise, especially for nuScenes, and we demonstrate in the experiments that Alligat0R is highly robust to this noise.
Figure 4: Distributions of overlap, scale ratio, and viewpoint angle in Cub3-all and Cub3-50 for nuScenes (left) and ScanNet (right).
Figure˜4 illustrates the distribution of our datasets across three geometric criteria introduced in RUBIK [loiseau2025rubik]: overlap percentage, scale ratio, and viewpoint angle. The Cub3-all dataset’s distribution is heavily skewed towards very challenging pairs, which are particularly difficult for cross-completion methods like CroCo. In contrast, our covisibility segmentation approach is designed to benefit from such challenging cases effectively.
Given our computational resources, Cub3 is currently limited to either autonomous driving scenarios from nuScenes, or indoor scenes from ScanNet. But such scenarios have very important applications, and given its size, variety, and challenges, Cub3 is already sufficient to validate Alligat0R—very much like the initial version of CroCo [weinzaepfel2022croco] was demonstrated as a "proof-of-concept":
  • Scale: With 5M annotated pairs, Cub3 provides sufficient data to train and validate our approach on both outdoor and indoor scenarios.
  • Geometric variety: It covers a wide range of overlaps, scales and viewpoint angles within the driving and indoor contexts.
@@
Our experiments on the RUBIK benchmark, which was specifically designed to evaluate challenging autonomous driving scenarios, as well as results on ScanNet1500 test set demonstrate the effectiveness of our approach. The promising results we obtain motivate future work to extend Cub3 to more diverse environments beyond urban driving scenes and indoor scenes. However, even within this specific domain, our method shows significant improvements over CroCo, particularly in geometrically challenging configurations.
@@ 5.1 Implementation Details
We implement Alligat0R using PyTorch and conduct all experiments on NVIDIA A100 GPUs. The model architecture follows the design described in Section˜3.1. The input images are classically resized to have a maximum dimension of 512 pixels while maintaining their aspect ratio. More information is provided in the supplementary material.
@@ 5.2 Main Results
Table 1: Results on RUBIK [loiseau2025rubik] and ScanNet1500 [dai2017scannet] for metric relative pose regression. For all experiments, fine-tuning is performed using Cub3-all.
RUBIK ScanNet1500 Pre-training Pre-Training Set Backbone $5^{\circ}$ / 0.5m $5^{\circ}$ / 2m $10^{\circ}$ / 5m $10^{\circ}$ / 0.25m $10^{\circ}$ / 0.5m $10^{\circ}$ / 1m CroCo Cub3-50 [Uncaptioned image] 4.4 21.8 48.1 48.3 64.1 69.3 CroCo Cub3-all [Uncaptioned image] 2.4 8.9 25.2 8.7 18.3 24.9 Alligat0R Cub3-50 [Uncaptioned image] 9.3 32.4 58.4 47.5 59.9 63.4 Alligat0R Cub3-all [Uncaptioned image] 24.0 55.3 82.3 78.2 90.1 92.5 CroCo Cub3-50 [Uncaptioned image] 12.4 38.3 66.7 75.7 87.4 91.5 CroCo Cub3-all [Uncaptioned image] 12.5 38.6 66.2 64.1 79.1 85.0 Alligat0R Cub3-50 [Uncaptioned image] 21.2 53.3 78.1 82.8 91.7 93.9 Alligat0R Cub3-all [Uncaptioned image] 24.6 60.3 81.9 85.5 92.5 95.1 From scratch Cub3-all [Uncaptioned image] 10.8 29.1 52.3 37.0 51.8 57.2
@@
Metric Relative Pose Regression Performance. We evaluate our proposed Alligat0R pre-training approach on the metric relative pose regression task and compare it with CroCo pre-training using the same architecture and fine-tuning protocol. Table˜1 presents the results on the RUBIK [loiseau2025rubik] benchmark, which is specifically designed to evaluate performance across different geometric challenges in autonomous driving scenarios, and on the indoor ScanNet1500 [dai2017scannet] benchmark.
@@
As shown in Table˜1, Alligat0R consistently outperforms CroCo across different training and evaluation configurations. Notably, when pre-trained on Cub3-all and fine-tuned with the backbone frozen, Alligat0R achieves significantly higher performance and reaches state-of-the-art on the RUBIK benchmark (55.3% success rate at $5^{\circ}$/2m and 82.3% at $10^{\circ}$/5m) compared to the CroCo model pre-trained on the same dataset (which only achieves 8.9% and 25.2% respectively in the frozen backbone configuration). This substantial improvement demonstrates the effectiveness of our covisibility segmentation pre-training approach, especially for challenging scenarios with varying degrees of overlap.
Impact of Training Data Distribution. A key observation is the contrasting behavior of Alligat0R and CroCo when trained on different data distributions. While CroCo performs similarly or better when trained on easier pairs (Cub3-50) compared to harder pairs (Cub3-all), Alligat0R shows a different trend. Our method systematically improves when trained on the full distribution (Cub3-all). This demonstrates that Alligat0R can effectively learn from challenging examples, while CroCo struggles to benefit from them.
@@
Figure 5: Performance of CroCo and Alligat0R across different geometric challenges on RUBIK. Results show accuracy at the $5^{\circ}$/2m threshold for models trained on different datasets (Cub3-50 or Cub3-all) and for frozen backbones. Alligat0R trained on Cub3-all consistently outperforms other configurations, particularly for challenging cases with low overlap, large scale differences, and extreme viewpoint changes.
@@
Figure˜5 further illustrates this advantage by breaking down performance across different overlap percentages, scale ratios and viewpoint angles. While both methods perform well on high-overlap pairs (80-100%), Alligat0R trained on Cub3-all maintains strong performance even as overlap decreases, whereas CroCo’s accuracy drops dramatically for pairs with less than 40% overlap. This demonstrates Alligat0R’s ability to generalize across a wide range of geometric configurations, which is crucial for real-world applications where overlap cannot be guaranteed.
@@
Fine-tuning Strategy Comparison. We also compare different fine-tuning strategies, specifically the impact of freezing versus unfreezing the backbone during fine-tuning. One can see that Alligat0R trained on Cub3-all is already quite high with a frozen backbone and always outperforms CroCo even when its backbone is unfrozen. This suggests that our pre-training strategy is better aligned with the downstream pose regression task.
@@
Comparison with Training from Scratch. To isolate the impact of our pre-training approach, we compare the performance of a model trained from scratch for relative pose regression versus those initialized with Alligat0R or CroCo pre-training. The bottom row of Table˜1 shows that training from scratch achieves notably lower performance, highlighting the value of the representations learned during pre-training.
Robustness to Label Noise. To address concerns about sensitivity to covisibility segmentation errors, we conducted a sensitivity analysis by training Alligat0R on ScanNet with 20% random label noise injected during pre-training. Surprisingly, this actually improved performance across all thresholds (+2.2%, +2.4%, +0.8% respectively), as shown in Table˜2. This counterintuitive result suggests that label noise acts as a form of regularization similar to label smoothing [szegedy2016rethinking], encouraging learning of more generalizable features. Importantly, this demonstrates that our approach remains robust even when covisibility annotations are imperfect.
Table 2: Sensitivity analysis: Impact of label noise on Alligat0R performance on ScanNet1500. Results show that 20% random label noise during pre-training actually improves performance, demonstrating robustness to annotation errors.
ScanNet1500
Method Noise $10^{\circ}$ / 0.25m $10^{\circ}$ / 0.5m $10^{\circ}$ / 1m
Alligat0R 0% 85.5 92.5 95.1
Alligat0R 20% 87.7 94.9 95.9
CroCo N/A 75.7 87.4 91.5
@@
Table 3: RUBIK Benchmark. The results are reported from [loiseau2025rubik]. Success rate (in %) for each method across individual geometric criterion bins. Best and second-best values for each column are shown in bold and underlined respectively.
@@
LoFTR [sun2021loftr] 87.2 88.4 47.2 17.5 5.0 51.6 10.1 2.3 0.6 43.2 27.9 15.1 0.0 24.9 185
ELoFTR [wang2024efficient] 56.4 90.3 50.8 22.1 6.3 51.2 15.6 4.4 0.7 42.2 30.8 18.2 0.1 26.6 124
ASpanFormer [chen2022aspanformer] 72.2 72.3 44.5 21.9 7.4 46.0 14.9 6.9 1.6 42.5 27.2 16.0 0.1 24.8 108
RoMa [edstedt2024roma] 67.0 98.3 84.5 52.7 20.2 71.2 43.2 26.6 8.3 57.5 56.2 44.1 3.0 47.3 614
DUSt3R [wang2024dust3r] 81.8 97.4 90.8 58.4 30.4 73.3 57.9 40.1 9.9 67.4 55.3 50.0 35.2 54.8 257
MASt3R [leroy2024grounding] 52.0 97.5 89.6 61.0 28.4 71.2 52.3 42.5 13.8 53.5 65.6 54.5 14.1 53.6 173
@@ 5.3 Qualitative Analysis
Covisibility and Reconstruction Visualization. Figure˜6 shows qualitative examples of the covisibility segmentation and CroCo reconstructions produced by our models on pairs from Cub3. Alligat0R successfully identifies covisible regions, occluded areas, and regions outside FOV across varying degrees of overlap and viewpoint changes. These visualizations provide insights into the geometric understanding acquired by our model. Conversely, while Croco succeeds in reconstructing masked covisible regions, the reconstructions are blurred in masked non-covisible regions as cross-completion is ill-posed in such areas.
 Reference    Target       Masked   CroCo recons.   Alligat0R segmentations
Figure 6: Qualitative comparison. CroCo’s reconstructions are blurred in masked non-covisible regions, while Alligat0R often correctly identifies covisible, occluded, outside FOV regions across varying degrees of overlap and viewpoint changes.
@@ 5.4 Comparison with state-of-the-art on RUBIK
In Table˜3, we evaluate the performance of Alligat0R fine-tuned for metric relative pose estimation against state-of-the-art methods on RUBIK [loiseau2025rubik]. Let us highlight that:
@@
This gives an advantage to Alligat0R over the other methods. However, RUBIK allows the evaluation over large ranges of three geometric criteria—overlap, scale ratio and viewpoint angle—which we found interesting. Despite this advantage, we argue that Alligat0R’s performance remains remarkable, as pretraining with CroCo, in contrast, leads to a significant performance drop (see Table˜1).
With an overall success rate of 60.3%, Alligat0R ranks first. Notably, it significantly outperforms all other methods on the hardest bin across the three geometric criteria, which confirms the advantage of its pretraining on difficult image pairs. While Alligat0R’s training set was highly skewed toward challenging pairs (see Figure˜4), one might expect a performance drop on easier bins. However, this does not happen: for instance, Alligat0R still ranks first for small scale ratios and small viewpoint angles. Finally, Alligat0R is among the fastest methods, as it directly regresses the relative pose rather than producing intermediate correspondences.
5.5 Out-of-Domain Generalization
To demonstrate the generalization capabilities of our method beyond the training domains, we evaluate Alligat0R on zero-shot correspondence estimation using the ETH3D dataset [schops2017multi]. We use the correlations from decoder features to estimate correspondences and measure performance using Average EndPoint Error (AEPE) as described in [an2025cross].
Zero-shot Correspondence Estimation. As detailed in the supplementary material, Alligat0R consistently outperforms CroCo when trained on the same datasets, and even surpasses CroCo v2 which was pre-trained on 5 datasets. This demonstrates that our covisibility segmentation pre-training learns more generalizable features for dense correspondence tasks.
Comparison with Official CroCo v2. We also compare against the official CroCo v2 model, fine-tuned on our Cub3-all dataset for pose regression. As shown in Table˜4, Alligat0R significantly outperforms CroCo v2 across all metrics, demonstrating the effectiveness of our covisibility segmentation approach even when compared to a model pre-trained on multiple datasets.
Table 4: Comparison with official CroCo v2 fine-tuned on Cub3-all dataset for pose regression on RUBIK benchmark. Alligat0R significantly outperforms CroCo v2 across all metrics.
RUBIK
Method $5^{\circ}$ / 0.5m $5^{\circ}$ / 2m $10^{\circ}$ / 5m
Alligat0R (Frozen) 24.0 55.3 82.3
CroCo v2 (Frozen) 5.8 15.8 34.1
Alligat0R (Unfrozen) 24.6 60.3 81.9
CroCo v2 (Unfrozen) 14.6 44.5 73.1
6 Limitations
While our approach demonstrates significant improvements over CroCo, it has several limitations:
  • First, our current evaluation is limited to urban driving scenes from nuScenes and indoor scenes from ScanNet. The performance of Alligat0R on more diverse environments, such as natural scenes or extreme weather conditions, remains to be investigated.
  • Second, our pre-training method requires ground truth depth maps and poses to produce dense covisibility annotations. While we investigated a fully automated pseudo ground truth generation for nuScenes, such a pre-processing is computationally intensive. This dependency on 3D data might limit the scalability of our approach to scenarios where such data is not readily available.
  • Finally, while we show promising results on metric relative pose regression, the effectiveness of our pre-training approach for other binocular vision tasks, such as stereo matching or optical flow estimation, remains to be explored.